# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

### v1

@@ Abstract
Canada experienced in 2023 one of the most severe wildfire seasons in recent history, causing damage across ecosystems, destroying communities, and emitting large quantities of $\text{CO}_{2}$. This extreme wildfire season is symptomatic of a climate-change-induced increase in the length and severity of the fire season that affects the boreal ecosystem. Therefore, it is critical to empower wildfire management in boreal communities with better mitigation solutions. Wildfire probability maps represent an important tool for understanding the likelihood of wildfire occurrence and the potential severity of future wildfires. The massive increase in the availability of Earth observation data has enabled the development of deep learning-based wildfire forecasting models, aiming at providing precise wildfire probability maps at different spatial and temporal scales. A main limitation of such methods is their reliance on coarse-resolution environmental drivers and satellite products, leading to wildfire occurrence prediction of reduced resolution, typically around $\sim 0.1$°. This paper presents a benchmark dataset: CanadaFireSat, and baseline methods for high-resolution: $100$ m wildfire forecasting across Canada, leveraging multi-modal data from high-resolution multi-spectral satellite images (Sentinel-2 L1C), mid-resolution satellite products (MODIS), and environmental factors (ERA5 reanalysis data). Our experiments consider two major deep learning architectures. We observe that using multi-modal temporal inputs outperforms single-modal temporal inputs across all metrics, achieving a peak performance of $60.3\%$ in F1 score for the 2023 wildfire season, a season never seen during model training. This demonstrates the potential of multi-modal deep learning models for wildfire forecasting at high-resolution and continental scale.
@@ 1 Introduction
As climate change accelerates, forests represent a key ecosystem as they act as one of the main terrestrial carbon sinks (Keenan and Williams, 2018), a shelter for a major part of Earth’s biodiversity (Lindenmayer and Franklin, 2013) during this new era of great extinction (Barnosky et al., 2011), and a critical environment for numerous fragile human communities (Fernández-Llamazares et al., 2021). Climate change endangers the role of forests as a carbon sink: it has been observed that several forest biomes are at risk of shifting from carbon sinks to carbon emitters (Pan et al., 2024). In particular, the boreal ecosystem is a subarctic biome in the high northern latitudes characterized by coniferous and mixed deciduous-coniferous forests. As the northernmost forest on the planet, boreal forests span $58\%$ of Canada’s land mass and are present across 7 other countries. To date, boreal forests represent one of the largest terrestrial carbon sinks, with approximately $367.3$ petagrams to $1715.8$ petagrams of carbon (Bradshaw and Warkentin, 2015). However, they are at risk of permafrost thaw due to land impacts (Li et al., 2021) and are increasingly subject to long and devastating wildfire seasons (McCarty et al., 2021).
@@
Wildfires greatly threaten many forest ecosystems due to the rapid shift in climate change-driven fire regimes. Indeed, despite a decreasing trend in the global burned area (mainly driven by a decrease of human-caused fires in African grasslands), (Bowman et al., 2020), areas burned by forest fires have seen a steady yearly increase of $\sim 5\%$ since 2001 (Tyukavina et al., 2022). In addition, forest fire events are burning longer and covering larger areas as extreme wildfire events’ frequency increases (Cunningham et al., 2024). As a consequence, wildfire $\text{CO}_{2}$ emissions reached $24\%$ of the total global emissions in 2023, corresponding to $8.6$ Gt of $\text{CO}_{2}$ (Jones et al., 2024; Liu et al., 2024). The observed amplification in wildfire severity is particularly true for the boreal ecosystem, representing roughly $70\%$ of the fire-related tree cover loss (Tyukavina et al., 2022) and where single wildfire events, like those during the 2023 Canadian wildfires season, can compete with annual emissions of major industrialized nations (Byrne et al., 2024). Locally, boreal wildfires have a direct impact on the land surface as they directly enhance permafrost thaw (Li et al., 2021; Zhao et al., 2024), and also contribute to vegetation shifts to more fire-prone grassland-/steppe-dominant landscapes, as well as dry peat (Zhao et al., 2024; McCarty et al., 2021). Those effects on local ecosystems, combined with wildfire $\text{CO}_{2}$ emissions, create a positive feedback mechanism fueling climate change. Finally, boreal wildfires emit large amounts of aerosols into the atmosphere, which can directly impact ice melt in the Arctic (Schmale et al., 2021) and are dangerous for public health, both locally and after smoke transportation thousands of kilometers away (Weilnhammer et al., 2021).
@@
The use of remote sensing data to map and monitor wildfires has expanded, with studies considering satellite-based observations of vegetative fuel conditions, individual fire events, and the impacts of smoke. Researchers have developed numerous wildfire tools, with three main use cases focused on the different phases of a wildfire: before a wildfire occurs (pre-fire), during a wildfire (active fire), and after (post-fire). For active fires, satellite products and models detect active fire ”hotspots” in near real-time (de Almeida Pereira et al., 2021; Růžička et al., 2022) or predict wildfire spread (Huot et al., 2022; Hoang et al., 2022), providing precious tools for wildfire management and decision-making. In the aftermath of wildfires (i.e, post-fire), methods to precisely segment the perimeter of burned areas (Rashkovetsky et al., 2021; Zhang et al., 2022) were developed to evaluate wildfires emissions, from $\text{CO}_{2}$ to black carbon, and estimate the impact of wildfires on the local ecosystem, natural resources, and communities. This study focuses on the first category (pre-fire) for wildfire prevention, shaped from a methodological perspective as a forecasting task. Wildfire forecasting, often referred to as wildfire susceptibility or likelihood modeling (Pelletier et al., 2023; Zhang et al., 2021), aims to predict the spatial probability of a wildfire occurring in a given time horizon and at a given spatial resolution (Prapas et al., 2022). This is done by producing wildfire probability maps. Wildfire forecasting is particularly useful for wildfire management by supporting staff and resource planning (Wotton, 2009).
@@
Wildfire forecasting is, by definition, a difficult task since it seeks to represent a complex and stochastic phenomenon. Fire susceptibility depends on several drivers: i) hydrometeorological conditions are the main variables that impact the suitability of vegetative fuels for combustion (ie, ’dryness’) and fire propagation (Krawchuk et al., 2009). Flammability is also directly linked to ii) the available biomass for combustion, which depends on iii) the type of vegetation and other indicators such as iv) dead or live fuel moisture (Krawchuk et al., 2009). Climate change makes those predictors for wildfire occurrence non-stationary. For example, meteorological patterns are highly variable in the boreal biome, which can lead to extreme fire seasons (McCarty et al., 2021). In addition, numerous vegetation changes have been observed or predicted, such as permafrost thaw (Li et al., 2021; Zhao et al., 2024), peatland destruction (Bourgeau-Chavez et al., 2022), or a shift from coniferous to deciduous forest (McCarty et al., 2021). Moreover, for effective wildfire forecasting, it is necessary to estimate v) the probability of ignition caused by humans or lightning (Pérez-Invernón et al., 2023) through proxies such as the proximity to human settlements. This variability implies that for similar environmental conditions, a wildfire may or may not occur based on latent variables for the model, which makes wildfire forecasting an especially complex task where ignition, even from lightning-only, is hard to model (Coughlan et al., 2021; Bates et al., 2021).
@@
Historically, wildfire forecasting was performed by producing fire weather indices, such as the Canadian Forest Fire Weather Index (FWI) or the National Fire Danger Rating System (NFDRS), which are mainly driven by meteorological and fuel moisture data. Fire weather indices aim to represent complex relationships between fire predictors through constraint equations based on assumptions that can lack robustness, such as focusing on specific forest types (e.g ”Pinus Banksiana”) or not taking into consideration the topography which can lead to necessary re-calibrations of the indexes for specific areas (Steinfeld et al., 2022; De Jong et al., 2016). Moreover, due to limitations in weather reanalysis or forecast data, fire weather indexes are often limited to short-term temporal horizons and coarse resolution. For instance, across Canada, the FWI has limitations in properly identifying the hydrometeorological conditions for combustion across all land cover types, and particularly so in peatlands (Waddington et al., 2012). Moreover, those indices cannot approximate the stochastic character of wildfire occurrence as they focus on flammability conditions. In parallel, traditional machine learning (ML) algorithms based on handcrafted features were proposed in (Martell et al., 1987, 1989) to identify drivers linked to wildfire occurrence, such as hydrometeorological conditions and human activities. These ML algorithms are limited in their ability to represent complex predictor relationships and possible spatio-temporal patterns, mostly because of the rigidity of the features used. Nevertheless, these methods are still widely used (Buch et al., 2023; Rodrigues et al., 2022) even in remote sensing (Pelletier et al., 2023; Chowdhury and Hassan, 2015).
@@
The growing availability of open remote sensing data (Reichstein et al., 2019; Camps-Valls et al., 2021), which enables the monitoring of large and remote regions, now allows mapping the drivers of fire susceptibility. This contributed to the emergence of wildfire forecasting models leveraging remote sensing imagery with traditional ML (Pelletier et al., 2023; Chowdhury and Hassan, 2015) and deep learning (DL) (Yang et al., 2021) methods. Different model designs have been used in the recent literature [^1] We will limit this review to neural network-based approaches, which are the closest to our proposition.: for one-dimensional inputs, the method of choice is the Multi-Layer Perceptron (MLP) (Buch et al., 2023; Bakke et al., 2023; Milanović et al., 2020), while for temporal-only inputs Long Short Term Memory (LSTM) networks have been proposed (Natekar et al., 2021) due to their ability to process sequences (e.g time series) by passing information along the sequence. When considering spatial inputs, convolutional neural networks (CNN) and vision transformer (ViT) have been explored (Prapas et al., 2022, 2023). Finally, for spatio-temporal data, architectures like convolutional LSTM (ConvLSTM) (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Bali et al., 2021), which join the sequence processing abilities of LSTM networks to the spatial awareness of CNNs, have been proposed. There is no clear consensus on the best model to use, as results seem to vary depending on the dataset characteristics (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Jain et al., 2020), region of interest, forecast horizon, and predictors list. In terms of data, most methods leverage hydrometeorological predictors, from reanalysis data like ERA5 or weather stations with spatial resolution varying from $\sim 27$ km to $4$ km. They also leverage remote sensing products linked to vegetation at higher resolution (up to $500$ m) and static factors linked to land cover and human activities (Kondylatos et al., 2022; Prapas et al., 2021; Bakke et al., 2023; Prapas et al., 2022; Bali et al., 2021). Those predictors are then re-sampled according to a target output resolution for the wildfire probability maps, varying from $0.25$° for global applications (Bakke et al., 2023; Prapas et al., 2022) to up to $1$ km for more localized regions (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021). For instance, in the context of large countries such as Canada, which spans thousands of kilometers, the coarse output resolution of current wildfire probability maps is $\sim 0.1$° (Bali et al., 2021). This represents an important limitation of current wildfire forecasting models and datasets, as such coarse wildfire probability maps prevent wildfire management from properly allocating resources globally and lead to the underestimation of potential smaller wildfires.
In this paper, we propose a large-scale, multi-modal, and spatio-temporal dataset to train high-resolution wildfire forecasting models and identify the best approach. First, we propose a benchmark dataset for high-resolution wildfire forecasting: CanadaFireSat [^2] Code for the data generation and the model benchmarking can be accessed, respectively, at github.com/eceo-epfl/CanadaFireSat-Data and github.com/eceo-epfl/CanadaFireSat-Model available on the HuggingFace Hub, which allows the forecasting of wildfire risk at a spatial resolution of $100$ m over the entirety of Canada. CanadaFireSat allows us to go beyond what was possible with previous datasets, such as (Huot et al., 2020) or (Prapas et al., 2021), both with a target resolution of $1$ km over the U.S. and the Eastern Mediterranean region, respectively. We enable high-resolution wildfire forecasting by resorting to temporal series of multi-spectral satellite images (multi-spectral satellite images have been shown to be predictive of wildfires (Yang et al., 2021; Pelletier et al., 2023)), which complement temporal series of environmental drivers from both reanalysis data (ERA5) and coarse resolution satellite products (MODIS) (Prapas et al., 2022), as shown in Figure 1. In the proposed CanadaFireSat, the role of multi-spectral images is to provide high-resolution information on local fire drivers thanks to the $10$ m resolution bands for Red, Green, Blue, and near infrared (NIR) of Sentinel-2. The environmental variables represent commonly-used fire drivers, from temperature to the normalized difference vegetation index (NDVI) at coarser multi-scale resolutions. Secondly, in this paper, we demonstrate the potential of learning multi-modal models for wildfire forecasting by benchmarking two state-of-the-art computer vision architectures on CanadaFireSat: ResNet (He et al., 2016) and ViT (Dosovitskiy, 2020). We consider three settings with varying inputs modality: [Uncaptioned image]satellite images only (Sentinel-2 at $10$m), [Uncaptioned image]environmental predictors only (ERA5 at $11$km, FWI at $0.25$°, MODIS at $1$km and $500$m), and [Uncaptioned image]satellite and environmental data, as represented in Figure 1.
@@
Figure 1: Representation of the CanadaFireSat benchmark and the high-resolution wildfire forecasting task.
Our results on CanadaFireSat demonstrate that deep learning models can outperform a knowledge-driven baseline (FWI), in both normal and extreme fire seasons by up to $+15\%$ in PRAUC and $+11\%$ in F1 score. Moreover, we showcase the importance of multi-modal approaches for enhanced performance, and that convolution-based models outperform transformer architectures across all setups on CanadaFireSat. Our best-performing model: ResNet-50 in setting [Uncaptioned image]satellite and environmental data, reaches a top performance of $51.2$ F1 score and $46.1$ PRAUC on the 2022 validation set and up to $60.3$ F1 score and $57.0$ in PRAUC for the extreme fire season year of 2023. In comparison, the best performing TeleVit model (Prapas et al., 2023) on the SeasFire global benchmark reaches $\sim 62\%$ PRAUC on their test set for the same 8-day horizon and at a resolution of $0.25$°, a resolution 270 times coarser than CanadaFireSat. Even if not directly comparable, these results showcase the great potential of high-resolution wildfire forecasting via multi-modal models.
@@
2 CanadaFireSat Dataset
In this section, we present CanadaFireSat, a benchmark dataset for high-resolution wildfire forecasting. First, we describe the sampling scheme for the selection of positive and negative data samples in Section 2.1. Then, in Section 2.2 we provide the set of predictors that we extracted and combined to build our multi-modal learning benchmark for high-resolution wildfire forecasting. Table 1 provides the key information on this dataset.

### v2

@@ Abstract
Canada experienced in 2023 one of the most severe wildfire seasons in recent history, causing damage across ecosystems, destroying communities, and emitting large quantities of $\text{CO}_{2}$. This extreme wildfire season is symptomatic of a climate-change-induced increase in length and severity of fire seasons affecting the boreal ecosystem. Therefore, it is critical to empower wildfire management in boreal communities with better monitoring solutions. Wildfire probability maps are an important tool for understanding the likelihood of wildfire occurrence and the potential severity of future wildfires. Fire forecasting tools based on Earth observation data exist, but they are limited both by the lack of label information and by their reliance on coarse-resolution environmental drivers and satellite products, which leads to wildfire occurrence prediction of reduced resolution, typically around $\sim 0.1$°. To tackle these two limitations, this paper presents a benchmark dataset, CanadaFireSat, and baseline methods for high-resolution wildfire forecasting at $100$ m across Canada. CanadaFireSat leverages multi-modal data from high-resolution multi-spectral satellite images (Sentinel-2), mid-resolution satellite products (MODIS), and environmental factors (ERA5). We experiment with convolutional (CNN) and transformer (ViT) architectures. We observe that using multi-modal temporal inputs outperforms single-modal temporal inputs across all metrics, achieving a peak performance of $60.3\%$ in F1 score for the 2023 wildfire season, a season never seen during model training. This demonstrates the potential of multi-modal deep learning for wildfire forecasting at high-resolution and continental scale.
@@ 1 Introduction
As climate change accelerates, forests represent a key ecosystem to protect as they act as one of the main terrestrial carbon sinks (Keenan and Williams, 2018), a shelter for a major part of Earth’s biodiversity (Lindenmayer and Franklin, 2013) , and a critical environment for numerous fragile human communities (Fernández-Llamazares et al., 2021). In particular, the boreal ecosystem is a subarctic biome in the high northern latitudes characterized by coniferous and mixed deciduous-coniferous forests. They represent one of the largest terrestrial carbon sinks, with approximately $367.3$ petagrams to $1715.8$ petagrams of carbon stored (Bradshaw and Warkentin, 2015). However, they are at risk of permafrost thaw due to land impacts (Li et al., 2021) and are increasingly subject to long and devastating wildfire seasons (McCarty et al., 2021).
@@
While wildfires severity is amplifying globally (areas burned by forest fires have seen a steady yearly increase of $\sim 5\%$ since 2001 (Tyukavina et al., 2022)), its effect is particularly devastating for the boreal ecosystem, representing roughly $70\%$ of the fire-related tree cover loss (Tyukavina et al., 2022) and where single wildfire events, like those during the 2023 Canadian wildfires season, can compete with annual $\text{CO}_{2}$ emissions of major industrialized nations (Byrne et al., 2024). Locally, boreal wildfires have a direct impact on the land surface as they directly increase permafrost thaw (Li et al., 2021; Zhao et al., 2024), and contribute to vegetation shifts to more fire-prone grassland-/steppe-dominant landscapes, as well as dry peat (Zhao et al., 2024; McCarty et al., 2021). Boreal forests span $58\%$ of Canada’s land mass: in this paper we use Canada as the area of interest to tackle the problem of wildfire forecasting in boreal ecosystems.
@@
The use of remote sensing data to map and monitor wildfires has expanded, with studies considering satellite-based observations of vegetative fuel conditions, individual fire events, and the impacts of smoke. Numerous wildfire tools exist, with three main use cases focused on the different phases of a wildfire: before a wildfire occurs (pre-fire), during a wildfire (active fire), and after (post-fire). For active fires, satellite products and models detect active fire ”hotspots” in near real-time (de Almeida Pereira et al., 2021; Růžička et al., 2022) or predict wildfire spread (Huot et al., 2022; Hoang et al., 2022), providing tools for wildfire management and decision-making. In the aftermath of wildfires (post-fire), methods to precisely segment the perimeter of burned areas (Hu et al., 2023; Zhang et al., 2022) were developed to evaluate wildfires emissions, from $\text{CO}_{2}$ to black carbon, and estimate the impact of wildfires on the local ecosystem, natural resources, and communities. This paper focuses on the pre-fire phase, shaped from a methodological perspective as a forecasting task. Wildfire forecasting, often referred to as wildfire susceptibility or likelihood modeling (Pelletier et al., 2023; Zhang et al., 2021), aims to predict the spatial probability of a wildfire occurring in a given time horizon and at a given spatial resolution. This is done by producing wildfire probability maps. Wildfire forecasting is particularly useful for wildfire management by supporting staff and resource planning (Wotton, 2009).
@@
Wildfire forecasting is, by definition, a difficult task since it seeks to represent a complex and stochastic phenomenon. Fire susceptibility depends on several drivers: i) hydrometeorological conditions are the main variables that impact the suitability of vegetative fuels for combustion (ie, ’dryness’) and fire propagation (Krawchuk et al., 2009). Fire susceptibility is also directly linked to ii) the available biomass for combustion, which depends on iii) the type of vegetation and other indicators such as iv) dead or live fuel moisture (Krawchuk et al., 2009). Climate change makes those predictors for wildfire occurrence non-stationary. For example, meteorological patterns are highly variable in the boreal biome, which can lead to extreme fire seasons (McCarty et al., 2021). In addition, numerous vegetation changes have been observed or predicted, such as permafrost thaw (Li et al., 2021; Zhao et al., 2024), peatland destruction (Bourgeau-Chavez et al., 2022), or a shift from coniferous to deciduous forest (McCarty et al., 2021). Moreover, for effective wildfire forecasting, it is necessary to estimate v) the probability of ignition caused by humans or lightning (Pérez-Invernón et al., 2023) through proxies such as the proximity to human settlements. This variability implies that for similar environmental conditions, a wildfire may or may not occur based on latent variables for the model, which makes wildfire forecasting an especially complex task where ignition, even from lightning-only, is hard to model (Coughlan et al., 2021; Bates et al., 2021).
@@
Historically, wildfire forecasting was performed by producing fire weather indices, such as the Canadian Forest Fire Weather Index (FWI) or the National Fire Danger Rating System (NFDRS), which are mainly driven by meteorological and fuel moisture data. Fire weather indices aim to represent complex relationships between fire predictors through constrained equations based on simplifying assumptions, such as the forest type (e.g ”Pinus Banksiana”) and neglecting the topography, leading to necessary recalibrations of the indices for specific areas (Steinfeld et al., 2022; De Jong et al., 2016). For instance, across Canada, the FWI has limitations in properly identifying the hydrometeorological conditions for combustion across all land cover types, and particularly so in peatlands (Waddington et al., 2012). Moreover, those indices cannot approximate the stochastic character of wildfire occurrence as they focus on flammability conditions. In parallel, traditional machine learning (ML) algorithms based on handcrafted features were proposed in (Martell et al., 1987, 1989) to identify drivers linked to wildfire occurrence, such as hydrometeorological conditions and human activities. These ML algorithms are limited in their ability to represent complex predictor relationships and possible spatio-temporal patterns, mostly because of the rigidity of the features used. Nevertheless, these methods are still widely utilized (Buch et al., 2023; Rodrigues et al., 2022) even in remote sensing (Maffei et al., 2021; Chowdhury and Hassan, 2015).
@@
The growing availability of open-access remote sensing data (Reichstein et al., 2019; Camps-Valls et al., 2021), which enables the monitoring of large and remote regions, now allows mapping the drivers of fire susceptibility. This accumulation of data contributed to the emergence of wildfire forecasting models leveraging remote sensing imagery with neural networks (Xu et al., 2025). When processing hydrometeorological data in the form of one-dimensional inputs (i.e. tabular data), the method of choice is the Multi-Layer Perceptron (MLP) (Buch et al., 2023; Bakke et al., 2023; Milanović et al., 2020), while for temporal series of tabular inputs, Long Short Term Memory (LSTM) networks have been proposed (Natekar et al., 2021) due to their ability to capture temporal relationships. When considering spatial inputs, convolutional neural networks (CNN) and vision transformer (ViT) have been explored (Prapas et al., 2022, 2023). Finally, for spatio-temporal data, architectures like convolutional LSTM (ConvLSTM) (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Bali et al., 2021), which join the sequence processing abilities of LSTM networks to the spatial awareness of CNNs, have been proposed. There is no clear consensus on the best model to use, as results seem to vary depending on the dataset characteristics (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Jain et al., 2020), region of interest, forecast horizon, and predictors list. In terms of data, most methods leverage hydrometeorological predictors, from reanalysis data like ERA5 or weather stations with spatial resolution varying from $\sim 27$ km to $4$ km. Finally, researchers resort to remote sensing products (characterizing the vegetation) at higher resolution (up to $500$ m) and static factors symptomatic of land cover type and human activities (Kondylatos et al., 2022; Prapas et al., 2021; Bakke et al., 2023; Prapas et al., 2022; Bali et al., 2021). The individual predictors are then re-sampled to the target resolution corresponding to the final wildfire probability map, varying from 0.250.25° for global applications (Bakke et al., 2023; Prapas et al., 2022) to up to $1$ km for localized regions (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021). For instance, in the context of large countries such as Canada, which spans thousands of kilometers, the output resolution of current wildfire probability maps is $\sim 0.1$° (Bali et al., 2021). This represents an important limitation, as such coarse wildfire probability maps prevent wildfire management from properly allocating resources at a finer scale and lead to the underestimation of potential smaller wildfires.
In this paper, we propose a multi-modal and spatio-temporal dataset covering Canada to enable high-resolution ($100$ m) wildfire forecasting and benchmark different models to demonstrate their potential. Our contributions are as follows:
  1. 1.
    We introduce a benchmark dataset, CanadaFireSat [^1] Code for the data generation and the model benchmarking can be accessed, respectively, at github.com/eceo-epfl/CanadaFireSat-Data and github.com/eceo-epfl/CanadaFireSat-Model, available on the HuggingFace Hub, for high-resolution wildfire forecasting at $100$ m over Canada in 8-day forecasting window. CanadaFireSat enables high-resolution wildfire forecasting by resorting to temporal series of multi-spectral images (Sentinel-2) complemented by temporal series of environmental drivers from both reanalysis data (ERA5) and coarse resolution satellite products (MODIS), as shown in Figure 1.
  2. 2.
    We investigate the impact of negative sampling on wildfire forecasting through the collection of two test sets across the 2023 extreme wildfire season for CanadaFireSat. Besides a classic test set following the same sampling strategy as the train and validation sets, where wildfire forecasting models show compelling performance, we also propose a hard test set sampled adversarially: this allows studying the lower-bound performance of models under extreme conditions, where ignition constitutes the key discriminating factor to identify potential wildfires.
  3. 3.
    We demonstrate the potential of learning multi-modal models for high-resolution wildfire forecasting by benchmarking two state-of-the-art computer vision architectures on CanadaFireSat: ResNet (He et al., 2016) and ViT (Dosovitskiy, 2020) across three settings with varying input modalities: [Uncaptioned image]satellite images only (Sentinel-2 at $10$ m), [Uncaptioned image]environmental predictors only (ERA5 at $11$ km, FWI at 0.250.25°, MODIS at $1$ km and $500$ m), and [Uncaptioned image]satellite and environmental data.
CanadaFireSat allows a big leap in terms of resolution with respect to what was possible with previous datasets, such as (Huot et al., 2020) or (Prapas et al., 2021), both using a target resolution of $1$ km over the U.S. and the Eastern Mediterranean region, respectively. Moreover, our results on CanadaFireSat demonstrate that: i) deep learning models outperform a knowledge-driven baseline (FWI) in both normal and extreme fire seasons, and ii) multi-spectral and hydrometeorological data complement each other, with multi-modal models providing the most accurate predictions.
@@
Figure 1: The CanadaFireSat benchmark and the high-resolution wildfire forecasting task.
@@
2 The CanadaFireSat Dataset
In this section, we present CanadaFireSat, a benchmark dataset for high-resolution wildfire forecasting. First, we describe the sampling scheme for the selection of positive and negative data samples in Section 2.1. Then, in Section 2.2 we detail the set of predictors extracted and combined to build our multi-modal learning benchmark for high-resolution wildfire forecasting. Table 1 summarizes CanadaFireSat’s main characteristics.